44. The height of a helicopter above the ground is givenby, where h is in meters and t is inseconds. After 2.00s, the helicopter releases a small mailbag. Howlong after its release does the mailbag reach the ground? Show allwork!
Answer #1
Height above ground, at t = 2.0 s
                  h = 3.00 t3 = 3.00x8 = 24.0 m
Velocity of the copter, after t = 2.0 s,
v = dh/dt = 9.0t2 = 9.0x2x2 = 36.0m/s
height h is increasing withtime. Velocity of the copter andhence that of the mailbag at the time of release are upward;positive.
The net displacement of the mailbag and acceleration ofgravity are downwaard and hence negative. a = g = - 9.8m/s2
          S = ut+ 0.5at2
        S = ut - 4.9t2
      -24 = 36.t - 4.9t2
    4.9.t2 - 36.t - 24 =0
     Solving, t = 7.96 s
Answer #2
First of all, you would need to find the height after 2.00seconds:
= 3.00(2.00)3 =3.00(8.00) = 24.0 m
Then you use a formula to find the final velocity before ithit the ground:
Vf2 = Vi2 +2gd
NOTE: Vf needs to be negative because it is goingdownward.
Vi = 0 m/s because it doesn't say the helicopter was movingup.
g = -9.81 m/s2
d = 24.0 m
-(Vf2 )= (0 m/s)2 + 2(-9.81m/s2)(24.0 m)
-(Vf2) = -470.88 or-471 m2/s2 as significantdigits
Then I can divid by negative to get the number to be positiveso I can find the square root of it.
√(471 m2/s2 ) = 21.7 m/s and makeit negative because it's going downward so its -21.7 m/s.
Now I can find the time finally.
Vf =Vi + gt
- 21.7 m/s = 0 m/s + (-9.81 m/s2)(t)
I divid both sides by -9.81 m/s2 and I get:
2.21 s = t
Answer #3
Hi!
Ok so starting with:
h = 3.00t3
Also they give us: t = 2.00s
So, plugging in t we get:
h = 3.00(2.00)3
h = 3.00(8.00)
h = 24.00 meters is when the mailbag is released!

What we have is:
distance = 24.00 meters
So we can use the formula: d = v0* t + (1/2)at2
where v0 = the initial velocity of the mailbag
   v = the finalvelocity of the mailbag
           a =acceleration acting on the mailbag
           t = time it takes to travel thatdistance

Since the mailbag is being released from the airplane,
there is no PUSH or initial velocity, its just being dropped.

So, we can say
v0 = 0
So, d = 0 + (1/2) at2
      24.00m = (1/2) (9.81m/s2)(t2)
      48.00m = (9.81m/s2)(t2)
      4.89s2 =(t2)                                                   sidenote:s2 is just a unit. When you                                                                                                divide 48m with 9.81m/s2 the                                                                                           m/s2 is the acceleration of                                                                                 gravity (meters per second per second)
    2.21s   =t
Hope this helps!